OpenStudy (anonymous):

A square and an isosceles triangle have the same perimeter. The sides of the square are x + 2 and the base of the triangle is x and sides x +5,find x.

OpenStudy (anonymous):



OpenStudy (anonymous):

x + x + 5 + x + 5 = x +2 + x +2 + x +2 + x +2

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\large P = 4(x+2) \] \[\large P = (x+5)+(x+5)+x \]

OpenStudy (anonymous):

um which is right i am confused

OpenStudy (anonymous):

all of them are right ways to do it

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

All of them are the *SAME* way to do it. ;)

OpenStudy (anonymous):

they arent answers, we set up the equation for you to solve for x

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

Yes.